# Perspective: Time irreversibility in systems observed at coarse resolution

### v2

@@ I Introduction
A broken time-reversal symmetry, or broken detailed balance, is essential for the existence of living matter [1, 2, 3]. Stated from a thermodynamic perspective, the emergence and persistence of life requires constant entropy production [4, 3]. These two statements are intuitive and may nowadays seem obvious, perhaps rightfully so. However, an attempt to unambiguously define time-reversal symmetry (and violations thereof) and consistently connect it to thermodynamics purely in terms of observations turns out to be rather non-trivial. Avoiding the conceptually much harder problem of the emergence of a positive entropy production from mechanics (classical or quantum) [5, 6], the main difficultly we are referring to in this perspective is to define, understand, and infer violations of detailed balance in a system’s dynamics when one cannot observe all degrees of freedom. There may be cases, where one at least has some knowledge about how the observable depends on the relevant (i.e., “slow”; to be made precise later), hidden microscopic degrees of freedom, and in some cases one does not even have such knowledge.
From a practical perspective, measurements on complex systems either have a finite resolution [7, 8, 9, 10] or do not resolve all relevant degrees of freedom—they only probe low-dimensional projections of the dynamics, such as the magnetization [11] or dielectric response [12], and observables such as molecular extensions or FRET efficiencies and lifetimes in single-molecule experiments [13, 14, 15, 16, 17, 18, 19, 20]. It is well known that these projections typically introduce memory [21, 22, 23]. In these situations, many microscopic states become “lumped” onto the same observable “state”. Similar projections emerge when introducing “reaction coordinates” to reduce the dimensionality of data to the essential observable(s) [24, 25, 26], which is taken a step further if one constructs kinetic (typically Markov) models with a small number of most relevant, long-lived states [27, 28, 29]. These projections and reductions significantly distort or mask signatures of a broken time-reversal symmetry in microscopic dynamics. The main aim of this perspective is not on biological applications [30], but to highlight why in defining and discussing time-reversal symmetry, and quantifying its violations, the precise underlying assumptions on the microscopic dynamics, the projection/coarse graining, and further reductions, are not merely a technical detail. These assumptions typically decide whether the conclusions that are drawn are physically correct or inconsistent, and in turn whether they lead to controversy.
@@ II Entropy production in Markovian systems coupled to an equilibrium bath
where in the last step we used Jensen’s inequality for the convex exponential function and Eq. (1), follows the detailed fluctuation theorem $\langle{\rm e}^{-\Delta s_{\rm inf}}\rangle=1$ and non-negativity of the entropy production (i.e., the second law) $\Delta S_{\rm inf}\geq 0$.
@@ II.1 Thermodynamic entropy production
While path-wise, stochastic definitions of the entropy production exist [34], throughout this perspective we focus on the deterministic, average values. We first address steady-state dynamics with invariant density $p_{\rm s}(x)$, i.e., we first consider strongly ergodic systems relaxing to an invariant density $p(x,\tau\to\infty)=p_{\rm s}(x)$, which requires $a(x)$ to be sufficiently confining. We either prepare the system in the steady state, i.e., $x_{\tau=0}\sim p_{\rm s}(x)$, or begin our observations after a sufficiently long time. Such systems may be in thermodynamic equilibrium (these are referred to as being in “detailed balance”), where no entropy is produced. Conversely, a system in the stationary state may also perform work $W$ against the friction [33]
@@ II.2 Equivalence of thermodynamic and informatic entropy production for overdamped Markov dynamics
where $\dot{x}_{\tau}\equiv\frac{\mathrm{d}x_{\tau}}{\rm d\tau}$ and note that the last term does not appear in the small-noise limit [43] but is neccessary in general (see Theorem 9.1 and Corollary in Chapter VI of Ref. [44]). Notably, it is well known that $\mathbbm{P}(|\dot{x}_{\tau}|<C)=0$ for any $C<\infty$ and any $\tau$, so Eq. (8) should cause (substantial) concern, since paths are locally almost surely non-differentiable. The issue may be resolved by recalling the support theorem of diffusion [45], according to which (loosely speaking) the probability measure of nowhere differentiable diffusion paths $(x_{\tau})_{0\leq\tau\leq t}$ has support on the closure of smooth paths $(y(\tau))_{0\leq\tau\leq t}\in C^{\infty}$, as long as said smooth approximations are interpreted according to the Stratonovich convention. The Stratonovich formulation is required for consistency between systems with “smooth” and “Brownian” inputs (see also [46]). In this context, we ought to make the replacement $x_{\tau}\to y(t)$ in Eq. (8) along the following lines.
@@ III.1 Markovian coarse-grained dynamics
Recall that as soon as the waiting time in the discrete states is not exponentially distributed or depends on previous states, the coarse-grained jump process is not Markovian. To make this explicit, only milestoning-type of coarse grainings may yield Markov jump dynamics in the presence of a time-scale separation, whereas lumping does not [55, 52, 53]. Beside not calling it “milestoning”, this fact was realized already by Kramers in his seminal work [58] (see Figure 2 and note that Kramers’ partitioning of configuration space is a prime example of milestoning; for more details see Supplemental Material in Ref. [54]). It is also very well established in the community building Markov-state models [55, 53]. Yet, it has only recently becoming appreciated in Stochastic Thermodynamics [52, 54, 57, 59, 60, 61] which we will address below in more detail.
Figure 2: Numerical evaluation of one-dimensional barrier crossing. (a, e) Dynamics $\mathrm{d}x_{\tau}=-\partial_{x}\phi(x)\mathrm{d}\tau+\sqrt{2}\mathrm{d}W_{\tau}$ evolving in a double-well potential $\phi(x)=x^{4}-3x^{2}+x/2$ simulated with $\mathrm{d}\tau=0.01$. The sample trajectory (a) is coarse-grained into two states by lumping $x\leq 0$ and $x>0$ (b) or milestoning with $A=-0.6$ and $B=0.6$ (f), where in the latter procedure the coarse-grained state changes once the milestones $A$ or $B$ are reached. Zoom into the initial part of the trajectory highlights frequent state changes (“fast re-crossings”) in the lumped dynamics (c) compared to (g) the milestoned dynamics. Waiting time density for state changes inferred as a histogram from a very long trajectory ($5\times 10^{7}$ time steps) for (d) lumped for the transition $x>0\to x\leq 0$ (Inset: for the transition $x\leq 0\to x>0$) and (h) milestoned for the transition $B\to A$ (Inset: for the transition $A\to B$) dynamics reveal that only the latter agrees with the desired exponential (black dashed line is $\lambda\exp(-\lambda t)$ for $\lambda^{-1}=$ mean waiting time). Note that the lumped trajectories (b, c) and waiting times (d) explicitly depend on the time resolution and do not converge as $\mathrm{d}\tau\to 0$.
@@ III.2 Non-Markovian coarse-grained dynamics
Before delving into specific strategies, we isolate two main points: (1) How can we reliably deduce that the underlying dynamics is out of equilibrium, i.e., whether $\Delta S_{\rm tot}>0$? (2) Can we infer a lower bound $\Delta S_{\rm bound}>0$ such that we know $\Delta S_{\rm tot}\geq\Delta S_{\rm bound}$? Given a coarse grained observation, one can generally of course not recover with certainty the full entropy production or an upper bound, since one may always miss out on parts of the system that are projected out.
@@ III.2.1 Lumped observables of continuous-space dynamics
If any of these observables is non-zero, this implies $\Delta S_{\rm inf}(q)>0$ and under the given assumptions therefore also $\Delta S_{\rm tot}(x)>0$, i.e., it implies, as desired, that the full dynamics is not at equilibrium. Of course, one always has to be careful when making deductions about the ensemble average from finite statistics, i.e., one has to make sure that the given averages truly deviate a statistically significant amount from 00.
@@
Disregarding waiting-time contributions (for reasons that will be detailed below) it is possible to estimate the steady-state entropy production $\dot{S}_{\rm tot}=\frac{\mathrm{d}}{\mathrm{d}t}\Delta S_{\rm tot}(x)$ from path measures of $(q_{\tau})_{0\leq\tau\leq t}$. If we let $\hat{\gamma}_{k}$ denote a particular set of $k+1$ consecutive observed states with $k\geq 1$, and $\mathbbm{P}(\hat{\gamma}_{k})$ the probability to observe this set along a (formally infinitely long) trajectory, then the $k$-th order estimator is given by [103]
$\dot{S}^{\rm est}_{k}\equiv\frac{1}{T_{k}}\sum_{\hat{\gamma}_{k}}\mathbbm{P}(\hat{\gamma}_{k})\ln\frac{\mathbbm{P}(\hat{\gamma}_{k})}{\mathbbm{P}(\theta\hat{\gamma}_{k})}\leq\dot{S}_{\rm tot}(x),$ (22)
where $T_{k}$ is the average duration of a trajectory $(q_{\tau})_{0\leq\tau\leq t}$ that visits $k+1$ distinct observed states (i.e., with $k$ observed state changes) and $\theta\hat{\gamma}_{k}$ is the corresponding time-reversed sequence of lumped $k+1$ states. If $(q_{\tau})_{0\leq\tau\leq t}$ is a $n$-th order semi-Markov process then $\dot{S}^{\rm est}_{n+k}=\dot{S}^{\rm est}_{n},\,\forall k>0$, and if $\hat{\Gamma}_{t}=\Gamma_{t}$ is a Markov process, we have $\dot{S}^{\rm est}_{1}=\dot{S}$. Note, however, that when $P(\hat{\gamma}_{k})$ are inferred from trajectory data, undersampling may cause deviations from these relations.
@@ III.2.2 Milestoned observables
First, note that, by definition, any milestoned trajectory lives on a discrete state space. Thus, trajectories can be exactly described by the sequence of states $q_{1},q_{2},\dots$ and the amount of time $t_{1},t_{2},\dots$ spent in the respective states. However, as generally the case for coarse grained dynamics, even if we decide to assume $(x_{\tau})_{0\leq\tau\leq t}$ to be Markovian, the coarse grained dynamics can have arbitrarily complicated memory and non-exponential and history-dependent waiting times. In order to see why milestoning with “good milestones” is beneficial, we recall the reasons for emergence of strong memory effects. First, by projecting onto localized cores, see Fig. 1, we inherently eliminate fast local recrossing events, and focus on long-time/larger-scale, yielding waiting times that are closer to being exponentially distributed. Second, if we select milestones to be small enough, we minimize the correlations between entry and exit points of milestones, and thereby reduce correlations in the sequence of visited states as well as correlations between waiting times in a given state and the sequence of preceding states. What is a “good” choice of milestones is a tricky question and in full generality is still an unsolved problem (see, e.g., Refs. [53, 55]). A “good” choice of milestones that are to yield a dynamics as close as possible to Markov will be the one for which microscopic trajectories will spend only short periods of time outside any milestone and correlations between sequences of milestones will be minimal [52].
@@
For lumped dynamics, we saw before that it is possible to infer lower bounds on the entropy production $\Delta S_{\rm tot}(x)$ by estimating $\Delta S_{\rm inf}(q)$. In particular, this could be done by separating $\Delta S_{\rm inf}(q)$ into a part obtained from the sequence of states and a part obtained from the times spent in each state before a transition into a certain direction [102]. Based on this rewriting, however, one can quite easily construct examples that, when applying the naive time reversal for milestoned dynamics $(q_{\tau})_{0\leq\tau\leq t}$, have $\Delta S_{\rm inf}(q)>0$ while $\Delta S_{\rm inf}(x)=\Delta S_{\rm tot}(x)=0$ [109] (note that, unlike for the lumped dynamics, the log-sum inequality is no longer helpful).
@@
This, at first surprising, result is caused by the non-commuting of time reversal and coarse graining, i.e., because for $(q_{\tau})_{0\leq\tau\leq t}=\mathcal{F}[(x_{\tau})_{0\leq\tau\leq t}]$ (where $\mathcal{F}$ denotes the milestoning functional) we may generally have $(q_{t-\tau})_{0\leq\tau\leq t}\neq\mathcal{F}[({x}_{t-\tau})_{0\leq\tau\leq t}]$. We stress that this result does not contradict the results in [110, 111], since a milestoned path weight $\mathbbm{P}((q_{\tau})_{0\leq\tau\leq t})$ is not a marginal path weight of the microscopic process $\mathbbm{P}((x_{\tau})_{0\leq\tau\leq t})$ [52]. The phenomenon was coined kinetic hysteresis and was addressed rigorously for first-order semi-Markov processes that arise as milestonings of thermodynamically consistent overdamped diffusions on a graph, including the Markovian limits [52]. Meanwhile, the results were extended to milestoning of Markov-jump dynamics [54] and milestonings of non-Markovian lumped dynamics [57].
@@
The fact that there are examples with $\Delta S_{\rm inf}(q)>0$ that have $\Delta S_{\rm inf}(x)=\Delta S_{\rm tot}(x)=0$ [109, 57], where in both cases, the backwards dynamics is naively defined to be the dynamics read backwards, has far-reaching consequences. Whenever the presence of milestoning cannot be excluded, even if we are allowed to assume that the microscopic dynamics $(x_{\tau})_{0\leq\tau\leq t}$ obeys steady-state dynamics as in Eqs. (2) or in Sec. III.1, we cannot bound the entropy production of $x$ by the time-reversal asymmetry of $(q_{\tau})_{0\leq\tau\leq t}$, i.e., by assessing directly $\Delta S_{\rm inf}(q)$. Therefore, all aforementioned inference strategies are prone to fail for milestoned processes, i.e., one can construct counterexamples (see Ref. [109]) to the inference strategies cited in Sec. III.2.1. Moreover, one can no longer deduce non-equilibrium from asymmetries in correlation functions of based on the observations $(x_{\tau})_{0\leq\tau\leq t}$, and approaches like thermodynamic uncertainty relations and speed limits do not necessarily apply, although future work might possibly extend these results to milestoned dynamics.
@@
How to estimate entropy production for milestoned dynamics? If we have further knowledge about the underlying microscopic dynamics, such as the existence of some Markov states or Markovian transitions, we may be able to make strict statements about the entropy production of the underlying dynamics [52, 59, 108]. Alternatively, if we happen to have access to the transition times (i.e., not only to the waiting times in the states but the time that the transitions between the states take), then we can infer non-equilibrium from violations of transition-path symmetries [80, 81, 82]. Without extra knowledge, but still assuming that $x$ is a steady-state overdamped Markov process, the presumably only (bulletproof) thing to do in general is to (given sufficient sampling) use higher-order estimators in Eq. (22) [103, 57] and intentionally disregard waiting times in order to avoid inconsistencies as shown in [52, 57, 109].

### v3

@@ I Introduction
A broken time-reversal symmetry, or broken detailed balance[1, 2, 3, 4, 5], is essential for the existence of living matter [6, 7, 8]. Stated from a thermodynamic perspective, the emergence and persistence of life requires constant entropy production [9, 8]. These two statements are intuitive and may nowadays seem obvious, perhaps rightfully so. However, an attempt to unambiguously define time-reversal symmetry (and violations thereof) and consistently connect it to thermodynamics purely in terms of observations turns out to be rather non-trivial. Avoiding the conceptually much harder problem of the emergence of a positive entropy production from mechanics (classical or quantum) [10, 11], the main difficultly we are referring to in this perspective is to define, understand, and infer violations of detailed balance in a system’s dynamics when one cannot observe all degrees of freedom. There may be cases, where one at least has some knowledge about how the observable depends on the relevant (i.e., “slow”; to be made precise later), hidden microscopic degrees of freedom, and in some cases one does not even have such knowledge.
From a practical perspective, measurements on complex systems either have a finite resolution [12, 13, 14, 15] or do not resolve all relevant degrees of freedom—they only probe low-dimensional projections of the dynamics, such as the magnetization [16] or dielectric response [17], and observables such as molecular extensions or Förster resonance energy transfer efficiencies and lifetimes in single-molecule experiments [18, 19, 20, 21, 22, 23, 24, 25]. It is well known that these projections typically introduce memory [26, 27, 28]. In these situations, many microscopic states become “lumped” onto the same observable “state”. Similar projections emerge when introducing “reaction coordinates” to reduce the dimensionality of data to the essential observable(s) [29, 30, 31], which is taken a step further if one constructs kinetic (typically Markov) models with a small number of most relevant, long-lived states [32, 33, 34]. These projections and reductions significantly distort or mask signatures of a broken time-reversal symmetry in microscopic dynamics. The main aim of this perspective is not on biological applications [35], but to highlight why in defining and discussing time-reversal symmetry, and quantifying its violations, the precise underlying assumptions on the microscopic dynamics, the projection/coarse graining, and further reductions, are not merely a technical detail. These assumptions typically decide whether the conclusions that are drawn are physically correct or inconsistent, and in turn whether they lead to controversy.
@@ II Entropy production in Markovian systems coupled to an equilibrium bath
where in the last step we used Jensen’s inequality for the convex exponential function and Eq. (1), follows the detailed fluctuation theorem $\langle{\rm e}^{-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Delta s_{\rm inf}}}\rangle=1$ and non-negativity of the entropy production (i.e., the second law) $\Delta S_{\rm inf}\geq 0$.
@@ II.1 Thermodynamic entropy production
While path-wise, stochastic definitions of the entropy production exist [39], throughout this perspective we focus on the deterministic, average values. We first address steady-state dynamics [47] with invariant density $p_{\rm s}(x)$, i.e., we first consider strongly ergodic systems relaxing to an invariant density $p(x,\tau\to\infty)=p_{\rm s}(x)$, which requires $a(x)$ to be sufficiently confining. We either prepare the system in the steady state, i.e., $x_{\tau=0}\sim p_{\rm s}(x)$, or begin our observations after a sufficiently long time. Such systems may be in thermodynamic equilibrium (these are referred to as being in “detailed balance”), where no entropy is produced. Conversely, a system in the stationary state may also perform work $W$ against the friction [38]
@@ II.2 Equivalence of thermodynamic and informatic entropy production for overdamped Markov dynamics
where $\dot{x}_{\tau}\equiv\frac{\mathrm{d}x_{\tau}}{\rm d\tau}$ and note that the last term does not appear in the small-noise limit [52] but is necessary in general (see Theorem 9.1 and Corollary in Chapter VI of Ref. [53]). Although it is often claimed in the literature that the last term only appears when interpreting the dynamics in the Stratonovich convention (see, e.g., Ref. [54]), this appears to be a pure misconception. First, the choice of convention is fixed when stating the equation of motion Eq. (2). Second, for additive noise ($D$ independent of $x$) the convention does not influence the statistics of paths $(x_{\tau})_{0\leq\tau\leq t}$ and consequently the action $A$ in Eq. (8) is also the same. In the Appendix, we summarize different strategies to derive Eq. (8), and highlight why for all strategies the only correct form of this equation has to include $\int_{0}^{t}\mathrm{d}\tau\nabla\cdot a(x_{\tau})/2$.
Notably, it is well known that $\mathbbm{P}(|\dot{x}_{\tau}|<C)=0$ for any $C<\infty$ and any $\tau$, so the $\dot{x}_{\tau}$ in Eq. (8) should cause (substantial) concern, since paths are locally almost surely non-differentiable. The issue may be resolved by recalling the support theorem of diffusion [55], according to which (loosely speaking) the probability measure of nowhere differentiable diffusion paths $(x_{\tau})_{0\leq\tau\leq t}$ has support on the closure of smooth paths $(y(\tau))_{0\leq\tau\leq t}\in C^{\infty}$, as long as said smooth approximations are interpreted according to the Stratonovich convention. The Stratonovich formulation is required for consistency between systems with “smooth” and “Brownian” inputs (see also [56]). In this context, we ought to make the replacement $x_{\tau}\to y(t)$ in Eq. (8) along the following lines.
@@ III.1 Markovian coarse-grained dynamics
Recall that as soon as the waiting time in the discrete states is not exponentially distributed or depends on previous states, the coarse-grained jump process is not Markovian. To make this explicit, only milestoning-type of coarse grainings may yield Markov jump dynamics in the presence of a time-scale separation, whereas lumping does not [65, 62, 63]. Beside not calling it “milestoning”, this fact was realized already by Kramers in his seminal work [68]. The Kramers result is displayed in Fig. 2 where the waiting-time density of a two-state system is shown after lumping (see Fig. 2b-d; far from the exponential line) and milestoning (see Fig. 2f-h; approaching the exponential in the limit of a very high and sharp barriers and well-separated milestones). It is also very well established in the community building Markov-state models [65, 63]. Yet, it has only recently becoming appreciated in Stochastic Thermodynamics [62, 64, 67, 69, 70, 71] which we will address below in more detail.
Figure 2: Numerical evaluation of one-dimensional barrier crossing. (a, e) Dynamics $\mathrm{d}x_{\tau}=-\partial_{x}\phi(x_{\tau})\mathrm{d}\tau+\sqrt{2}\mathrm{d}W_{\tau}$ evolving in a double-well potential $\phi(x)=x^{4}-3x^{2}+x/2$ simulated with $\mathrm{d}\tau=0.01$. The sample trajectory (a) is coarse-grained into two states by lumping $x\leq 0$ and $x>0$ (b) or milestoning with $A=-0.6$ and $B=0.6$ (f), where in the latter procedure the coarse-grained state changes once the milestones $A$ or $B$ are reached. Zoom into the initial part of the trajectory highlights frequent state changes (“fast re-crossings”) in the lumped dynamics (c) compared to (g) the milestoned dynamics. Waiting time density for state changes inferred as a histogram from a very long trajectory ($5\times 10^{7}$ time steps) for (d) lumped for the transition $x>0\to x\leq 0$ (Inset: for the transition $x\leq 0\to x>0$) and (h) milestoned for the transition $B\to A$ (Inset: for the transition $A\to B$) dynamics reveal that only the latter agrees with the desired exponential (black dashed line is $\lambda\exp(-\lambda t)$ for $\lambda^{-1}=$ mean waiting time). Note that the lumped trajectories (b, c) explicitly depend on the time resolution and do not converge as $\mathrm{d}\tau\to 0$ (more precisely, in the mathematical idealization $\mathrm{d}x_{\tau}=-\partial_{x}\phi(x_{\tau})\mathrm{d}\tau+\sqrt{2}\mathrm{d}W_{\tau}$, there will be a diverging number of recrossings such that the waiting time density (d) for a fixed trajectory of finite length $t$ converges to a delta function at 00).
@@ III.2 Non-Markovian coarse-grained dynamics
Before delving into specific strategies, we isolate two main points: (1) How can we reliably deduce that the underlying dynamics are out of equilibrium, i.e., whether $\Delta S_{\rm tot}>0$? (2) Can we infer a lower bound $\Delta S_{\rm bound}>0$ such that we know $\Delta S_{\rm tot}\geq\Delta S_{\rm bound}$? Given a coarse grained observation, one can generally of course not recover with certainty the full entropy production or an upper bound, since one may always miss out on parts of the system that are projected out.
@@ III.2.1 Lumped observables of continuous-space dynamics
If any of these observables is non-zero, this implies $\Delta S_{\rm inf}(q)>0$ and under the given assumptions therefore also $\Delta S_{\rm tot}(x)>0$, i.e., it implies, as desired, that the full dynamics are not at equilibrium. Of course, one always has to be careful when making deductions about the ensemble average from finite statistics, i.e., one has to make sure that the given averages truly deviate a statistically significant amount from 00.
@@
Disregarding waiting-time contributions (for reasons that will be detailed below) it is possible to estimate the steady-state entropy production $\dot{S}_{\rm tot}=\frac{\mathrm{d}}{\mathrm{d}t}\Delta S_{\rm tot}(x)$ from path measures of $(q_{\tau})_{0\leq\tau\leq t}$. If we let $\hat{\gamma}^{(k_{i}:k_{j})}_{k}$ denote a subsequence $k_{i},k_{i+1},\ldots k_{j}$ of a particular set of $k+1$ consecutive observed states with final state $\hat{\gamma}^{({\rm f})}_{k}$ for $k\geq 1$, and $\mathbbm{P}[\hat{\gamma}^{(k_{i}:k_{j})}_{k}]$ the probability to observe this (sub)sequence along a (formally infinitely long) trajectory, then the $k$-th order estimator is given by [113]
$\dot{S}^{\rm est}_{k}\equiv\frac{1}{T}\sum_{\hat{\gamma}_{k}}\mathbbm{P}[\hat{\gamma}^{(1:k+1)}_{k}]\ln\frac{\mathbbm{P}[\hat{\gamma}^{(\rm f)}_{k}|\hat{\gamma}^{(1:k)}_{k}]}{\mathbbm{P}[(\theta\hat{\gamma}_{k})^{(\rm f)}|(\theta\hat{\gamma}_{k})^{(1:k)}]},$ (22)
where $T$ is the average waiting time per state in the observed trajectories and $\theta\hat{\gamma}_{k}$ is the corresponding time-reversed sequence of lumped $k+1$ states. If $(q_{\tau})_{0\leq\tau\leq t}$ is a $n$-th order semi-Markov process then $\dot{S}^{\rm est}_{n+k}=\dot{S}^{\rm est}_{n},\,\forall k>0$ and $\dot{S}^{\rm est}_{n}\leq\dot{S}_{\rm tot}$ [113]. If $\hat{\Gamma}_{t}=\Gamma_{t}$ is a Markov process, we have $\dot{S}^{\rm est}_{1}=\dot{S}$. Note, however, that when $P(\hat{\gamma}_{k})$ are inferred from trajectory data, undersampling may cause deviations from these relations.
@@ III.2.2 Milestoned observables
First, note that, by definition, any milestoned trajectory lives on a discrete state space. Thus, trajectories can be exactly described by the sequence of states $q_{1},q_{2},\dots$ and the amount of time $t_{1},t_{2},\dots$ spent in the respective states. However, as generally the case for coarse grained dynamics, even if we decide to assume $(x_{\tau})_{0\leq\tau\leq t}$ to be Markovian, the coarse grained dynamics can have arbitrarily complicated memory and non-exponential and history-dependent waiting times. In order to see why milestoning with “good milestones” is beneficial, we recall the reasons for emergence of strong memory effects. First, by projecting onto localized cores, see Fig. 1, we inherently eliminate fast local recrossing events, and focus on long-time/larger-scale, yielding waiting times that are closer to being exponentially distributed. Second, if we select milestones to be small enough, we minimize the correlations between entry and exit points of milestones, and thereby reduce correlations in the sequence of visited states as well as correlations between waiting times in a given state and the sequence of preceding states. What is a “good” choice of milestones is a tricky question and in full generality is still an unsolved problem (see, e.g., Refs. [63, 65]). A “good” choice of milestones that are to yield a dynamics as close as possible to Markov will be the one for which microscopic trajectories will spend only short periods of time outside any milestone and correlations between sequences of milestones will be minimal [62]. Note that the (somewhat counterintuitive) concurrent need for small milestones and short times spent outside the milestones for the emergence of Markovian dynamics highlights that approaching such a Markovian limit does by no means exist in general, but instead is a very special case.
@@
For lumped dynamics, we saw before that it is possible to infer lower bounds on the entropy production $\Delta S_{\rm tot}(x)$ by estimating $\Delta S_{\rm inf}(q)$. In particular, this could be done by separating $\Delta S_{\rm inf}(q)$ into a part obtained from the sequence of states and a part obtained from the times spent in each state before a transition into a certain direction [112]. Based on this rewriting, however, one can quite easily construct examples that, when applying the naive time reversal for milestoned dynamics $(q_{\tau})_{0\leq\tau\leq t}$, have $\Delta S_{\rm inf}(q)>0$ while $\Delta S_{\rm inf}(x)=\Delta S_{\rm tot}(x)=0$ (note that, unlike for the lumped dynamics, the log-sum inequality is no longer helpful; for a concrete example of how erroneous entropy production estimates emerge when applying the naive time reversal, see Ref. [119]).
@@
This, at first surprising, result is caused by the non-commuting of time reversal and coarse graining, i.e., because for $(q_{\tau})_{0\leq\tau\leq t}=\mathcal{F}[(x_{\tau})_{0\leq\tau\leq t}]$ (where $\mathcal{F}$ denotes the milestoning functional) we may generally have $(q_{t-\tau})_{0\leq\tau\leq t}\neq\mathcal{F}[({x}_{t-\tau})_{0\leq\tau\leq t}]$ (note that this statement equally holds for underlying dynamics $(x_{\tau})_{0\leq\tau\leq t}$ that represent overdamped or underdamped diffusion or Markov jump processes on discrete state-space). We stress that this result does not contradict the results in [120, 121], since a milestoned path weight $\mathbbm{P}((q_{\tau})_{0\leq\tau\leq t})$ is not a marginal path weight of the microscopic process $\mathbbm{P}((x_{\tau})_{0\leq\tau\leq t})$ [62]. The phenomenon was coined kinetic hysteresis and was addressed rigorously for first-order semi-Markov processes that arise as milestonings of thermodynamically consistent overdamped diffusions on a graph, including the Markovian limits [62]. Meanwhile, the results were extended to milestoning of Markov-jump dynamics [64] and milestonings of non-Markovian lumped dynamics [67].
@@
The fact that there are examples with $\Delta S_{\rm inf}(q)>0$ that have $\Delta S_{\rm inf}(x)=\Delta S_{\rm tot}(x)=0$ [119, 67], where in both cases, the backwards dynamics is naively defined to be the dynamics read backwards, has far-reaching consequences. Whenever the presence of milestoning cannot be excluded (in particular, in all theoretical works claiming to address a general class of experiments), even if we are allowed to assume that the microscopic dynamics $(x_{\tau})_{0\leq\tau\leq t}$ obeys steady-state dynamics as in Eqs. (2) or in Sec. III.1, we cannot bound the entropy production of $x$ by the time-reversal asymmetry of $(q_{\tau})_{0\leq\tau\leq t}$, i.e., by assessing directly $\Delta S_{\rm inf}(q)$. Therefore, all aforementioned inference strategies are prone to fail for milestoned processes, i.e., one can construct counterexamples (see Ref. [119]) to the inference strategies cited in Sec. III.2.1. Moreover, one can no longer deduce non-equilibrium from asymmetries in correlation functions of based on the observations $(x_{\tau})_{0\leq\tau\leq t}$, and approaches like thermodynamic uncertainty relations and speed limits do not necessarily apply, although future work might possibly extend these results to milestoned dynamics.
@@
How to estimate entropy production for milestoned dynamics? If we have further knowledge about the underlying microscopic dynamics, such as the existence of some Markov states or Markovian transitions, we may be able to make strict statements about the entropy production of the underlying dynamics [62, 69, 118]. Useful information abut the microscopic dissipation is also encoded in transition-path times, i.e. the duration of successful transition, corresponding to the timespan between the last exit of the microscopic trajectory from a milestone to the first entrance into a new one. If we indeed have access to these transition times, then we can infer non-equilibrium from violations of transition-path symmetries [90, 91, 92]. Without extra knowledge, but still assuming that $x$ is a steady-state overdamped Markov process, the presumably only (bulletproof) thing to do in general is to use saturated-order estimators in Eq. (22) [113] and intentionally disregard waiting times in order to avoid inconsistencies as shown in [62, 67, 119].